I need to have you show me how to calculate the answer to:
0.4r-1.2=0.3r+0.6
next step 4r-12=3r+6
next step 4r-12+-12=3r+6+-12
next step 4r=3r+-6
next step 6+4r=3r+-6+6
next step 10r=3r
next step 10/10r=3r/10
next step r=0.3

Answer by Earlsdon(5024)

You started out fine, but went a little astray on the way down.
0.4r - 1.2 = 0.3r + 0.6 Multiply through by 10 to clear the decimals.
4r - 12 = 3r + 6 Add 12 to both sides.
4r = 3r + 18 Subtract 3r from both sides.
r = 18
Check:
0.4(18) - 1.2 = 0.3(18) + 0.6
7.2 - 1.2 = 5.4 + 0.6
6 = 6 The solution (r = 18) is correct.